OpenStudy (anonymous):

Simplify: -15h – 6j + 8h – j

OpenStudy (cwrw238):

add like terms eg -15h+ 8h = -7h

OpenStudy (unklerhaukus):

\[-15h – 6j + 8h – j\]\[=-15h + 8h – 6j – j\]\[=(-15+ 8)h +( – 6 – 1)j\]\[=\]